OpenStudy (anonymous):

Again guys ?? <3 |5x + 2| = 17

OpenStudy (turingtest):

how did it get solved last time? 5x+2=17 5x+2=-17 now you can take it from here

OpenStudy (anonymous):

3,-19/5

OpenStudy (anonymous):

x=3 or -3

OpenStudy (anonymous):

no x=3 because 5(3)=15 +2 =17